# Paying to Do Better: Games with Payments between Learning Agents

### v1

@@ Abstract
In repeated games, such as auctions, players typically use learning algorithms to choose their actions. The use of such autonomous learning agents has become widespread on online platforms. In this paper, we explore the impact of players incorporating monetary transfers into their agents’ algorithms, aiming to incentivize behavior in their favor. Our focus is on understanding when players have incentives to make use of monetary transfers, how these payments affect learning dynamics, and what the implications are for welfare and its distribution among the players. We propose a simple game-theoretic model to capture such scenarios. Our results on general games show that in a broad class of games, players benefit from letting their learning agents make payments to other learners during the game dynamics, and that in many cases, this kind of behavior improves welfare for all players. Our results on first- and second-price auctions show that in equilibria of the “payment policy game,” the agents’ dynamics can reach strong collusive outcomes with low revenue for the auctioneer. These results highlight a challenge for mechanism design in systems where automated learning agents can benefit from interacting with their peers outside the boundaries of the mechanism.
@@ 1 Introduction
Autonomous learning agents have become widespread on online platforms, playing an ever-larger role in many markets and economic ecosystems. One prominent example is the billion-dollar market of online ad auctions run by internet giants like Google, Meta, and Microsoft, in which, due to the speed of these auctions, bidding traffic is controlled by various automated bidding agents, provided either by the platforms themselves, or by a third-party. Typically, these agents get some high-level instructions from their users about their objectives and allowed action space, and then they interact with other agents in long sequences of repeated games (which could include thousands or millions of auctions per hour), using some learning algorithm to optimize the long-term payoffs for their users.
@@
We are interested in repeated games of this sort where players use learning agents to play on their behalf. We take as a point of departure the observation that, in this setting, users may allow their agents to make payments to other agents in order to affect the game dynamic. In this paper, we study the effect of considering such payments between the agents on the long-term outcomes of the interaction.
@@
The idea that payments and other types of financial benefits outside of the mechanism can affect behavior in the mechanism has been observed before, in areas ranging from blockchain fee markets, where it has become one of the guiding principles in the evolving design of these markets [^1] While research in this area has focused on manipulations involving the auctioneer (the “miner” who allocates space for transaction to be included in the next “block”), we study the scenario where payments take place among the bidding agents themselves. (see [24, 79] and [8, 9, 25]), to fair division [64, 65]. We defer a more extended discussion of related work to Appendix A.
@@
The possibility that learning agents could transfer payments among themselves during their game dynamics raises several basic questions. First, when do players have incentives to let their agents use payments? Second, how do payments between the agents affect learning dynamics? And third, what are the long-term implications of payments among the agents on the players’ utilities, the social welfare, and in auctions, the revenue for the auctioneer?
@@
As in much of the prior literature (e.g., [16, 28, 32, 38, 46, 53, 90]), we assume agents are no-regret learners; that is, they use learning strategies that result in outcomes satisfying the no-regret condition in the long term. Unfortunately, as is well known, in many games no-regret agents can end up playing strategies that yield low utility compared to cooperative outcomes that could potentially have been obtained in the game. E.g., in the prisoner’s dilemma, they end up defecting, since that is the dominant strategy. This has led to a great deal of interest in designing mechanisms where we can get, in some sense, better outcomes. Notably, in contrast to works in mechanism design in which an external entity provides incentives or recommendations aiming to improve welfare [7, 10, 62, 88], in our scenario there is no external entity, and players are only interested in their own individual payoffs.
@@
Before giving an example, let us sketch the general structure of the model we have in mind. As mentioned, there is an underlying game and the model of what we call a payment policy game (or just a payment game) proceeds in two phases. In the first phase, each player $i$ chooses a payment policy, that determines, for each other agent $j$, how much agent $i$ will pay agent $j$ at each round in the agents’ dynamics as a function of the outcome (i.e., the joint action profile performed by the agents) in the underlying game in that round. In the second phase, the agents play the underlying game repeatedly, choosing their actions according to their learning algorithm (taking the payments into account). See Section 2 for further details.
@@
Warm-up example: With this background, consider the variant of prisoner’s dilemma [^2] Note that this variant is somewhat nonstandard, in that the payoffs are not symmetric. shown in Figure 1. As mentioned above, standard analysis shows that when both players use no-regret agents, each agent will quickly learn to play its dominant strategy, which is to defect (i.e., play strategy $D$); the dynamics will converge to the Nash equilibrium of the stage game where both agents always defect and each player gets an expected payoff of $\nicefrac{{1}}{{3}}$.
@@
Now we ask what can player $1$ achieve if he allows his agent to make payments to the other agent during the game dynamics? Consider the payment policy where player $1$’s agent pays player $2$’s agent $\nicefrac{{1}}{{3}}+\epsilon$ if player $2$’s agent cooperates, for some $\epsilon>0$, and pays nothing if player 2’s agent defects. It is easy to see that the game with these payments is dominance solvable: by repeatedly eliminating strictly dominated strategies, we are left only with the strategy $(D,C)$ as the unique equilibrium, and so the ensuing dynamics will converge to that outcome (see Appendix B), with payoffs $(2-\nicefrac{{1}}{{3}}-\epsilon,\nicefrac{{1}}{{3}}+\epsilon)$ for the two players.
It turns out that this outcome can be obtained as an ($\epsilon$-)equilibrium [^3] In an $\epsilon$-equilibrium, no player can increase their payoff by more than $\epsilon$ by altering their strategy. of the payment game. There is also a second $\epsilon$-equilibrium of the payment game, obtained by switching the roles of players $1$ and $2$, but due to the asymmetry in the payoffs of the underlying game, the social welfare in this second equilibrium is $1$, not $2$. Importantly, however, both equilibria are Pareto improvements over the Nash equilibrium of the underlying game, and the players’ welfare significantly increases compared to the game without payments.
We show, using a similar analysis, that we get these equilibria and improvement in social welfare in general prisoner’s dilemma games whenever the maximum welfare gap in the game is large (above a factor of $2$). Moreover, while in the game without payments the price of anarchy ($PoA$) and price of stability ($PoS$) —the ratio between the social welfare obtained from the worst-case (resp., best-case) Nash equilibrium and the optimal social welfare—are unbounded, with payments, both $PoS$ and $PoA$ are bounded by a factor of $2$ in the symmetric case, and in the asymmetric case $PoS$ is bounded, but $PoA$ is not, due to the welfare gap between the two equilibria of the payment game. See Appendix B.
@@
Our results: In this paper, we initiate the study of the incentives of players to use payments with their learning agents and the potential impact of such payments. We analyze first- and second-price auctions as our main case study in Section 3. In the following sections, we study general properties of payment games; Section 4 focuses on unilateral payments and Section 5 focuses on two-player games. Our results in all these settings provide comparative statics between the outcomes of learning dynamics without payments and those obtained in equilibria of the payment game.
More specifically, in second-price auctions, Theorem 1 shows that players have incentives to provide payment policies to their agents, and that in equilibrium of the payment game, the players may reach, in the long-term limit, full collusion where they capture almost the entire welfare, leaving the auctioneer with vanishing revenue. In first-price auctions, Theorem 2 shows that also here players have incentives to use payments, and equilibria of the payment game lead to low (but still positive) revenue. [^4] En route, we also provide the first simple dynamic approaching the minimum-revenue coarse correlated equilibrium of the first-price auction shown in [35]; see Observation 1. Then, in Section 4 (Theorems 3, and 4), we show that in a very broad class of finite games, [^5] Finite games are games with discrete action sets and bounded payoffs. Note that in the general model of auctions, the bid space is continuous, and so our analysis in this setting is different. there is at least some player who has the incentive to use payments, and we characterize cases where a payment policy by a single agent can lead the dynamics to the optimal welfare outcome while improving the payoff for the associated player as well. In Section 5, we show that in two-player games, payments always lead to Pareto improvements compared to the game without payments (Theorem 5), and that in a broad class of games, even including games with strictly dominant strategies (as our prisoner’s dilemma example shown above), at least one of the players would strictly benefit from using payments, and so not using payments is not stable (Theorem 6).
@@
From a higher-level perspective, our results highlight a challenge for auction design and market design more broadly: we show that users of learning agents quite generally have incentives to allow their automated agents to make payments during their interactions. When these incentives are coupled with the right technology, particularly, sophisticated AI agents and flexible transaction media such as those on blockchain platforms, one can easily imagine an outlook where agents are trading among themselves “under the hood,” and the associated markets change their behavior, equilibria, and outcomes. Our results demonstrate that these changes can be very significant, underlining, on the one hand, the potential improvement in efficiency, but on the other hand, the need to better understand these changes for concrete markets, to be able to design them accordingly.
@@ 2 Model
In our model, we consider scenarios where we have players using learning agents in some repeated game, and study the setting where players can augment their agents by allowing them to make payments to other agents that depend on the actions of these agents during the game dynamics. The learning algorithms used by the agents are assumed to satisfy the regret-minimization property, [^6] We provide the definition of regret minimization and other standard definitions in Appendix C. but agents are not restricted to using any particular algorithm, and the players themselves are interested in their long-term payoffs. As our analysis relies on the regret-minimization property itself and considers the outcomes in the limit $T\rightarrow\infty$, where $T$ is the number of game rounds, our results apply regardless of the specific regret bounds of the learning algorithms used, as long as the agents achieve regret sublinear in $T$. Our model has the following components.
@@
  • Payment policies: Each player $i$ chooses a policy that determines, for each action profile $s\in S$ and player $j\in[n]\setminus\{i\}$, the payment $p_{ij}(s)\in\mathbb{R}_{+}$ from agent $i$ to agent $j$ when the outcome in the most recent step is $s$. We assume that payments are bounded: $p_{ij}(s)\in[0,M]$, where $M\in\mathbb{R}_{+}$ is some large constant.
@@
  • Since the analysis of long-term payoffs ignores finite-time costs, we assume w.l.o.g. that if a player is asymptotically (in the limit $T\rightarrow\infty$) indifferent between making a payment and not making the payment, the player prefers not making the payment.
The model above defines a payment-policy game (in short, a payment game) between users of learning agents that is specified by $\Gamma$, $A$, and $T$. In order to understand the basic properties of payment-policy games and how the incentives of players change compared to games without payments, our analysis of equilibria and potential deviations focuses on one-shot, full-information payment-policy games.
A fundamental question that arises in our context is: “When would players have incentives to use payment policies?” Conversely, what are the conditions for a game to be “stable” in the sense that players do not have incentives to use such manipulations?

### v2

@@ Abstract
In repeated games, such as auctions, players typically use learning algorithms to choose their actions. The use of such autonomous learning agents has become widespread on online platforms. In this paper, we explore the impact of players incorporating monetary transfer policies into their agents’ algorithms, aiming to influence behavior in their favor through the dynamics between the agents. Our focus is on understanding when players have incentives to make use of monetary transfers, how such payments may affect learning dynamics, and what the implications are for welfare and its distribution among the players. We propose a simple and general game-theoretic model to capture such scenarios. Our results on general games show that in a very broad class of games, self-interested players benefit from letting their learning agents make payments to other learners during the game dynamics, and that in many cases, this kind of behavior improves welfare for all players. Our results on first- and second-price auctions show that in equilibria of the “payment policy game,” the agents’ dynamics reach strong collusive outcomes with low revenue for the auctioneer. These results raise new questions and highlight a challenge for mechanism design in systems where automated learning agents can benefit from interacting with their peers in the digital ecosystem and outside the boundaries of the mechanism.
@@ 1 Introduction
Autonomous learning agents have become widespread on online platforms, playing an increasingly pivotal role in many markets and economic ecosystems. A prominent example is the multi-billion-dollar online industry of ad auctions, operated by major companies such as Google, Meta, Amazon, and Microsoft. Due, in part, to the speed and complexity of these auctions, automated bidding has emerged as the dominant approach, with bidding traffic managed by various automated agents, provided either directly by the platforms themselves, or by third-party platforms. Typically, these agents get some high-level instructions from their users about their objectives and allowed action space, and then they interact with other agents in long sequences of repeated games (which could include thousands or millions of auctions per hour), using some learning algorithm to optimize the long-term payoffs for their users.
@@
We are interested in repeated interactions of this sort where players use learning agents to play a repeated game on their behalf. We take as a point of departure the observation that, in this setting, users may wish to allow their agents to make payments to other agents in order to influence the course of the game’s dynamic in their favor. In this study, we model and analyze the effect of considering such payments between the agents on the long-term outcomes of the interaction.
@@
The idea that payments and other types of financial benefits outside of a mechanism can affect behavior in the mechanism has received considerable attention in areas ranging from blockchain fee markets, where it has become one of the guiding principles in the evolving design of these markets [^1] While research in this area has focused on manipulations involving the auctioneer (the “miner” who allocates space for transaction in the next “block”), we study the scenario where payments take place among the bidding agents themselves. (see [28, 94] and [11, 12, 29]), to fair division [78, 80]. We defer a more extended discussion of related work to Section 3.
@@
The possibility that learning agents could transfer payments among themselves during their game dynamics raises several basic questions. First, when do players have incentives to let their agents use payments? Second, how do payments between the learning agents affect their dynamics? And third, what are the long-term implications of payments among the agents on the players’ utilities and the social welfare? (And in the application of auctions, the revenue for the auctioneer?)
@@
As in much of the prior literature (e.g., [19, 32, 38, 46, 56, 63, 105]), we model agents as no-regret learners; that is, they use learning strategies that result in outcomes satisfying the no-regret condition in the long term. Unfortunately, as is well known, in many games no-regret agents can end up playing strategies that yield low utility compared to cooperative outcomes that could potentially have been obtained in the game. E.g., in the prisoner’s dilemma, they end up defecting, since that is the dominant strategy. This has led to a great deal of interest in designing mechanisms where we can get, in some sense, better outcomes. Notably, unlike the literature on mechanism design, where a platform or some external entity provides incentives or recommendations aiming to improve welfare [10, 13, 76, 103], in our case, there is no external entity. All payments are made at the expense of the players, who are solely interested in maximizing their own individual payoffs.
@@
Before giving an example, let us sketch the general structure of the model we have in mind. As mentioned, there is an underlying game, and the model of what we call a payment policy game (or just a payment game) proceeds in two phases. In the first phase, each player $i$ chooses a payment policy, that determines, for each other agent $j$, how much agent $i$ will pay agent $j$ at each round in the agents’ dynamics as a function of the outcome (i.e., the joint action profile performed by the agents) in the underlying game in that round. In the second phase, the agents play the underlying game repeatedly, choosing their actions according to their learning algorithm (taking the payments into account). See Section 2 for the formal description of the model.
@@
Warm-up example: With this background, consider the variant of the prisoner’s dilemma [^2] Note that this variant is somewhat nonstandard, in that the payoffs are not symmetric. game shown in Figure 1. As mentioned above, standard analysis shows that when both players use no-regret learning agents, each agent will quickly learn to play its dominant strategy, which is to defect (i.e., play strategy $D$); the dynamics will converge to the Nash equilibrium of the stage game where both agents always defect and each player gets an expected payoff of $\nicefrac{{1}}{{3}}$.
@@
Now we ask what can player $1$ achieve if she allows her agent to make payments to the other agent during the game dynamic? Consider the payment policy where player $1$’s agent pays player $2$’s agent $\nicefrac{{1}}{{3}}+\epsilon$ if player $2$’s agent cooperates, for some $\epsilon>0$, and pays nothing if player 2’s agent defects. It is easy to see that the game with these payments is dominance solvable: by repeatedly eliminating strictly dominated strategies, we are left only with the strategy $(D,C)$ as the unique equilibrium, and so the ensuing dynamics will converge to that outcome (see Appendix A), with payoffs $(2-\nicefrac{{1}}{{3}}-\epsilon,\nicefrac{{1}}{{3}}+\epsilon)$ for the two players.
It turns out that this outcome can be obtained as an ($\epsilon$-)equilibrium [^3] In an $\epsilon$-equilibrium, no player can increase their payoff by more than $\epsilon$ by altering their strategy. of the payment game. There is also a second similar $\epsilon$-equilibrium of the payment game, obtained by switching the roles of players $1$ and $2$, but due to the asymmetry in the payoffs of the underlying game, the social welfare in this second equilibrium is $1$, not $2$. Importantly, however, both equilibria are Pareto improvements over the Nash equilibrium of the underlying game, and the players’ welfare significantly increases compared to the game without payments (note, though, that welfare is not necessarily optimal).
We show, using a similar analysis, that we get these equilibria and improvement in social welfare in general prisoner’s dilemma games whenever the maximum welfare gap in the game is large (above a factor of $2$). Moreover, while in the game without payments the price of anarchy ($PoA$) and price of stability ($PoS$)—the ratio between the social welfare obtained from the worst-case (respectively, best-case) Nash equilibrium and the optimal social welfare—are unbounded, with payments, both $PoS$ and $PoA$ are bounded by a factor of $2$ in the symmetric case, and in the asymmetric case $PoS$ is bounded, but $PoA$ is not, due to the welfare gap between the two equilibria of the payment game. See Appendix A for more details.
@@
Our contribution: We study the incentives of players to use payments with their learning agents and the potential impact of such payments among learning agents on learning dynamics and equilibrium outcomes.
@@
From a high-level perspective, our results highlight a challenge for auction design and market design more broadly: we show that users of learning agents quite generally have incentives to allow their automated agents to make payments during their interactions. When these incentives are coupled with the right technology—particularly, sophisticated AI agents and flexible transaction media, such as those on blockchain platforms—one can easily imagine an outlook where agents trade among themselves “under the hood,” and the associated markets change their behavior, equilibria, and outcomes. Our results demonstrate that these changes can be very significant, underlining, on the one hand, a potential improvement in efficiency, but on the other hand, the need to better understand the potential impact on concrete systems, to be able to design them accordingly.
We propose a model of these interactions for general games in Section 2. As our main case study, we analyze first- and second-price auctions in Section 4. In the following sections, we study general properties of payment games: Section 5 focuses on the incentives players have to initiate unilateral payments, and Section 6 focuses on the special case of two-player games. Our results in all these settings provide comparative statics between the outcomes of learning dynamics without payments and those obtained with learning agents in equilibria of the payment game between their users.
Summary of main theorems: In second-price auctions, Theorem 1 shows that players have incentives to provide payment policies to their agents (i.e., using learning agents in the auction without using payments is not a Nash equilibrium), and that in an equilibrium of the payment game, in the long run, the players may reach full collusion, where they capture almost the entire welfare, leaving the auctioneer with vanishing revenue as $T\rightarrow\infty$. This result applies for any number of players with generic (i.e., not exactly equal) valuations.
In first-price auctions, Theorem 2 shows that also here, players have incentives to use payments. So using automated bidders without using payments is not an equilibrium. This first part of the theorem extends to any number of players (see Appendix B for further details). For the special case of two players, the theorem shows that equilibria of the payment game lead to low (but in this case, still positive) revenue. [^4] En route, we present what is, to the best of our knowledge, the first simple dynamic that approaches the minimum-revenue coarse correlated equilibrium of the first-price auction as established in [41]; see Observation 1.
In Section 5, we show that under very broad conditions, players have incentives to use payments with their learning agents, also beyond auctions. Theorems 3 and 4, show that in a broad class of finite games, [^5] Finite games are games with discrete action sets and bounded payoffs. Note that in the general model of auctions, the bid space is continuous, and so our analysis in that setting is distinct from that of Theorems 3 and 4. there is at least some player who has the incentive to use payments, and we characterize cases where a payment policy applied by a single agent can drive the dynamics toward the optimal-welfare outcome while improving the payoff for the associated player as well.
In Section 6, we analyze the important case of bilateral interactions. Two-player games have unique properties in payment games, intuitively, stemming from the fact that any payment must be desirable to both players, as there are no other externalities on the associated agents from other interactions. Theorem 5 shows that in two-player games, payments always result in Pareto improvements compared to the game without payments. Furthermore, Theorem 6 demonstrates that in a broad class of games—extending even to those with strictly dominant strategies, as illustrated in our prisoner’s dilemma example above—at least one player will strictly benefit from using payments. Consequently, not using payments is not a stable behavior.
@@ 2 Model
We consider scenarios where we have players using learning agents in some repeated game, and study the setting where players can augment their agents by allowing them to make payments to other agents based on the actions of these agents during the game dynamics. The learning algorithms used by the agents are assumed to satisfy the regret-minimization property, [^6] For completeness, we provide standard definitions in Appendix C. but agents are not restricted to using any particular algorithm, and the players themselves are interested in their own long-term payoffs. As our analysis relies on the regret-minimization property itself and considers outcomes in the limit $T\rightarrow\infty$, where $T$ is the number of game rounds, our results apply regardless of the specific regret bounds or details of the learning algorithms used, as long as the agents achieve regret sublinear in $T$. Our model has the following components.
@@
  • Payment policies: Each player $i$ chooses a policy for her own agent that determines, for each action profile $s\in S$ and player $j\in[n]\setminus\{i\}$, the payment $p_{ij}(s)\in\mathbb{R}_{+}$ from agent $i$ to agent $j$ when the outcome in the most recent step is $s$. We assume that payments are bounded: $p_{ij}(s)\in[0,M]$, where $M\in\mathbb{R}_{+}$ is some large constant.
@@
  • Since in the analysis of long-term payoffs, finite-time costs vanish, we assume w.l.o.g. that if a player is asymptotically (in the limit $T\rightarrow\infty$) indifferent between making a payment and not making the payment, the player prefers not making the payment.
The model above defines a payment-policy game (in short, a payment game) between users of learning agents that is specified by $\Gamma$, $A$, and $T$. In order to understand the basic properties of payment-policy games and how the incentives of players change compared to games without payments, our analysis of equilibria and potential deviations focuses on one-shot, full-information payment-policy games. A fundamental question that arises is: when would players have incentives to use payment policies? Conversely: what are the conditions for a game to be “stable” in the sense that players do not have incentives to use such manipulations?
@@
3 Further Related Work
Learning in games: This paper follows a long research tradition on learning and dynamics in games, from early work of Brown, Robinson, Blackwell, and Hannan in the 1950s [16, 23, 52, 91], through seminal work in the following decades [47, 48, 54] and continuously since [1, 18, 30, 31, 45, 56, 61, 62, 63, 64, 66, 74, 98]. See [27, 55] for an overview of the foundations of the field. The notion of regret has been central throughout this work as a tool to define learning outcomes and objectives and design algorithms to achieve them.
Regret minimization in auctions: No-regret learning has been widely studied in auctions, with two prominent lines of work focusing on the price of anarchy in various auction formats with regret-minimizing bidders [17, 25, 32, 56, 92, 95, 99] and on the problem of inferring bidder preferences, where such bidder behavior is either assumed or empirically verified [49, 82, 83, 84]. Other work has studied automated bidding algorithms for auctions with more complex features, such as budget constraints [14, 34, 45, 44, 70], or learning reserve prices [26, 75, 96]. In [3, 87], regret minimization is used as a prediction model for bidder behavior. Importantly, [82, 87] provide empirical evidence from large auction datasets showing that bidder behavior in practice is, by and large, consistent with no-regret learning outcomes. Learning dynamics in first- and second-price auctions and their convergence properties have been studied in a broad range of work [15, 21, 32, 33, 43, 45, 44, 63, 64, 70].
Metagames: A recent line of work, following [63, 64], studies metagames between users of learning agents, in which players set parameters for their agents, which then repeatedly interact in some underlying game (e.g., an auction, a market, or an abstract normal-form game). The utilities in the metagame are determined by the long-term outcomes of the agents’ dynamics. These papers focus on analyzing the average outcomes of regret-minimizing agents and their implications on users’ incentives to misreport their preferences to their own agents. Further research has explored metagames in randomized auctions [2] as well as strategic budget selection in auctions [4, 42, 73] and its implications to welfare and revenue.
Our model is related to the metagame framework but addresses a distinct interaction type that has not been analyzed for learning agents: payments between agents. We show that payments between agents have a markedly different effect on the agents’ dynamics, leading to different analyses and results in auctions and other games compared to the strategic parameter reports studied in prior work. For example, in both first- and second-price auctions, equilibrium bids and revenues are lower in our payment-based scenario than in prior metagame results [63, 2].
Optimization against learning agents: An different line of research on interactions with learners considers scenarios where an optimizer interacts with a learning agent, such as an auctioneer optimizing auction rules against a learning buyer [22]. This has been extended to studying cases with multiple buyers [24], general two-player games [6, 35], Bayesian games [71], and repeated contracting [51]. While that line of work shares the premise of players committing to using learning algorithms, its focus is different: it does not study payments between the agents, and furthermore, it addresses an optimization problem where a single player solves for the best response to a fixed learning algorithm, while we study a game (where none of the players is a Stackelberg leader) and the equilibrium outcomes that arise from the long-term dynamics of multiple learning agents.
Cooperation in repeated games: The question of how distributed and self-interested players can achieve cooperative or efficient outcomes is a broad topic, studied across various, often separate, fields. A classic approach addresses this challenge through mechanism design [81, 85] and implementation theory [57, 72]. Somewhat closer to our work are papers that study how an external mechanism designer can minimize the exogenous payments she needs to make to the players in order to implement particular outcomes in the game (see, e.g., [10, 76, 103]). By contrast, we study the impact of payments that automated agents could make among themselves; our analysis does not consider a mechanism-design problem [^7] Designing payment protocols to facilitate particular types of payments between the players to optimize social welfare is an interesting research avenue, but we do not pursue it here. and there are no external funds or a directing hand. An additional strand of work from the reinforcement learning (RL) and artificial intelligence literature studies how distributed RL algorithms can jointly reach cooperative outcomes in sequential social dilemmas [9, 39, 60, 69, 102]. A third line of work considers the notion of program equilibrium [68, 88, 89, 100], in which each player declares a program and each program can read the commitments made in the other programs and condition actions in the game on those commitments. This model, however, is very different from our model.
In comparison to research on collaborative outcomes from the perspective of cooperative game theory [79], we note that while in our payment game the players do manage to share, to some extent, the welfare they obtain, the game is not cooperative. In particular, an equilibrium of the payment game is not necessarily an element of perhaps the most standard solution concept in cooperative games, the core [50]. This can be clearly seen, for instance, in our analysis of first-price auctions, where in equilibrium the bidders do not play an action profile in the core of their game (thinking of the auction mechanism as set, and the game as taking place only between the bidders).
Games with monetary transfers: The concept of using payments to incentivize desired behaviors in games is a well-established idea. Payments between players have been studied in the context of social choice [53] and fair division [78, 80], and are a central concept in the broad literature on contract theory and principal-agent problems. For an introductory overview on this literature, we refer interested readers to [20, 36, 67].
Most work in this area involves a “principal” who provides incentives to agents, thereby facing a contract-design or optimization problem rather than acting as a strategic player in the game. Notable papers on contracts more closely related include [58], which studies multi-lateral payment contracts in one-shot interactions; [51, 104], which examine learning in repeated contracts; and [90], which explores scenarios where a player extracts fees from others by using binding contracts to alter the equilibrium of the game. Our model however, diverges significantly from the standard contract framework. Instead of focusing on principal-agent dynamics or external incentives injected into the game, we analyze scenarios where players choose to augment their learning agents with the capacity to make payments. Our primary interest lies in how these payments influence the agents’ learning dynamics and, consequently, the outcomes of the games they play. Other work on games with payments related to ours include [37] and [65], who study sequential games with transfers in which players make sequential commitments, responding rationally to the commitments made by previous players. They study conditions where the sequential commitment structure leads to a unique subgame perfect equilibrium with full welfare. These works do not deal with learning agents and their setting and analysis differ significantly from ours.